# Learning to Communicate using Contrastive Learning

### v1

@@ 5.6 Protocol Representation Probing
Figure 6: DBSCAN (Ester et al., 1996) clustering results of messages produced by CACL after reduced in dimensions using t-SNE (Van der Maaten and Hinton, 2008). Exemplary clusters are shown with their corresponding observational patterns. Specifically, two clusters correspond to messages sent when the goal is visible and another agent is visible respectively. The other two clusters of when only individual agents are visible.
@@
Similar to Lin et al. (2021), we cluster messages generated from 10 evaluation episodes to qualitatively assess how informative CACL’s protocol is. The messages are first compressed to a dimension of 2 using t-SNE (Van der Maaten and Hinton, 2008) and then clustered using DBSCAN (Ester et al., 1996). We look at each cluster’s messages and their corresponding observations to extract any patterns and semantics captured. As shown in Figure 6, we observe a cluster of messages for observations when the goal is visible and another one when another agent is visible. Two clusters correspond to agents seeing neither the goal nor another agent. Notably, the messages in these clusters can come from different agents in different episodes, demonstrating that agents can indeed communicate symmetrically. The clusters indicate that CACL learns to compress meaningful, global state information in messages, allowing other agents to reasonably learn this semantic information.
@@
Table 2 shows the classification results for the two probing tests. Goal visibility is an easier task and all methods’ messages can be effectively used to determine it. In the more difficult goal location task, all methods perform above chance (20%) but CACL’s protocol significantly outperforms baselines. Contrastive learning across different agents’ messages can enable CACL to learn a more global understanding of location from their egocentric viewpoint. By encoding the goal’s spatial information, CACL agents are more likely able to move directly towards it, and reduce episode length. If other methods simply communicate that a goal is found, agents know to alter their search but are not as precise in direction. This explains why AEComm, PL, and DIAL perform better than IAC but worse than CACL, which also learns much quicker as shown in Figure 2. For completeness, we also provide similar classification results with a one-layer (linear) probe in Appendix A.5
@@ 7 Conclusion and Future Work
This work introduces an alternative perspective for learning to communicate in decentralized MARL based on the relationship between sent and received messages within a trajectory. Drawing inspiration from multi-view learning, we ground communication using contrastive learning by considering agents’ messages to be encoded views of the same state. We empirically show that our method leads to better performance through a more consistent, common language and learns to communicate more global state information. We believe this work solidifies contrastive learning as an effective perspective for learning to communicate and hope it invigorates research into contrastive methods for communication with a focus on consistency. Furthermore, by establishing the connection between multi-view SSL, which has traditionally focused on images, and communication in MARL, we hope to encourage more cross-domain research. Finally, we see contrastive learning as a potential method for simulating human language evolution, and hope to inspire research in this direction.
@@ Appendix A Appendix
A.1 Environment Details
@@
Figure 7: Visual illustration of the environments used. Left: Predator-Prey, taken from Koul [2019]. Middle: Find-Goal, taken from Lin et al. [2021]. Right: Traffic-Junction, taken from Singh et al. [2018]
A.1.1 Predator-Prey
We modify the Predator-Prey implementation by Koul [2019]. Our Predator-Prey has a higher communication and coordination requirement than the original Predator-Prey environment. Specifically, for a prey to be captured, it has to be entirely surrounded (i.e. the prey cannot move to another grid position in any actions).
@@ A.1.1 Predator-Prey
A.1.2 Find-Goal
We use the Find-Goal environment implementation provided by Lin et al. [2021]. The agents have the goal to find where the goal is in a 15x15 grid world with obstacles.
Unlike in Lin et al. [2021], each agent has a 3x3 field of view (instead of 7x7) to make the task more difficult. Each agent receives a reward of 1 for reaching a goal and an additional reward of 5 if all agents reach the goal. We use a step penalty of -0.01 and an obstackle density of 0.15.
@@ A.1.2 Find-Goal
A.1.3 Traffic-Junction
We use the Traffic-Junction environment implementation provided by Singh et al. [2018]. The gridworld is 8x8 with 1 traffic junction. The rate of cars being added has a minimum and maximum of 0.1 and 0.3. We use the easy version with two arrival points and 5 agents. Agents are heavily penalized if a collision happens and have only two actions, namely gas and brake.
@@ A.1.3 Traffic-Junction
A.2 Architecture and Hyperparameters
Figure 8: Architectural illustration for algorithms with communication. To remove communication, the message head is disabled. Grounding module is only relevant to CACL and AEComm. The former is a loss function and the latter is a decoder to reconstruct the encoded observation.
Figure 8 illustrates the components of the architecture used in this work, similar to [Lin et al., 2021]. A message head is only used for algorithms with communication, namely CACL, AEComm, PL and DIAL. The Grounding Module refers to mechanisms to ground the messages produced by the message head, used in CACL and AEComm. Unless specified otherwise, we fix all hidden layers to be a size of 32.
We experimented with using the output of the GRU, or hidden state, to condition the message head. Empirically we found that directly conditioning on the observation encoding, as in Lin et al. [2021], lead to more stable learning dynamics.
The observation encoder output values of size 32. For Predator-Prey and Traffic Junction, a one-layer fully-connected neural network is used as observation encoder. For Find-Goal, same as Lin et al. [2021], we use a two-layer convolutional neural network followed by a 3-layer fully-connected neural network.
For the message encoder, it outputs values of size 8 in Predator-Prey and Find-Goal with one hidden layer. It outputs values of size 16 in Traffic-Junction with two hidden layers. These configurations are selected based on the best performance of the baseline communication learning algorithm used - DIAL. Messages received are concatenated before passing to message encoders. For all the methods with communication, they produce messages of length 4 ($D=4$) with a sigmoid function as activation. All models are trained with the Adam optimizer [Kingma and Ba, 2014].
@@ A.2 Architecture and Hyperparameters
A.3 Predator-Prey Capture Rate
@@
A.4 Positive Listening
This section describes the loss function we implemented for positive listening, based on Eccles et al. [2019]. Given two policies $\pi^{i}$ and $\overline{\pi^{i}}$ of agent $i$ where the latter is the policy with messages zeroed out in the observations, and a trajectory $\tau$ of length $T$, the positive listening loss is written as:
@@ A.4 Positive Listening
A.5 Protocol Representation Probing: 1-Layer
@@
A.6 Zero-shot Cross-play
@@
An advanced form of coordination is working with partners you have not seen during training [Hu et al., 2020]. Previous work has focused on coordination through actions [Carroll et al., 2019, Lupu et al., 2021] or pre-test grounding with a common dataset [Gupta et al., 2021] but to our knowledge, no previous work has succeeded in learning a linguistic communication protocol that is robust to zero-shot partners. To assess this advanced robustness, we take trained agents from different methods and random seeds and evaluate them with each other (i.e., zero-shot cross-play) in Predator-Prey and Find-Goal. Given two communication learning methods, $m_{1}$ and $m_{2}$, we sample two agents from each method for Predator-Prey and for Find-Goal, we average over sampling two agents from one method and one agent from the other and vice-versa. For intra-method cross-play, $m_{1}=m_{2}$, we evaluate agents that were trained with the same method but from different random seeds, so they have not been trained with each other. For inter-method cross-play, $m_{1}\neq m_{2}$, we sample agents from two different methods and pair them with each other. Each pairing is evaluated for 10 random seeds each with 10 evaluation episodes. Given that agents are trained in self-play [Tesauro, 1994] without regard for cross-play, we expect severe performance dips.
@@
A.7 Broader Impact
More multi-agent learning systems will be deployed in the real-world as further progress is made in fields of multi-agent learning like MARL. We expect communication to play an essential role in these systems given how real-world problems are inherently complex and partially observable in most cases. Our focus on the decentralized communication setting contribute to the capability of learning more effective and consistent communication protocols. Having more consistent protocols improve mutual intelligibility and pave the way to multi-agent systems in which agents can communicate with unseen agents or even humans.

### v3

@@ 5.6 Protocol Representation Probing
Figure 6: DBSCAN (Ester et al., 1996) clustering results of messages produced by CACL after dimensionality reduction with t-SNE. We show four semantically meaningful clusters with their corresponding labels: messages sent when the goal is visible, when another agent is visible, and two clusters correspond to only the individual agents visible in the observation.
@@
Similar to Lin et al. (2021), we cluster messages generated from 10 evaluation episodes to qualitatively assess how informative CACL’s protocol is. The messages are first compressed to a dimension of 2 using t-SNE (Van der Maaten & Hinton, 2008) and then clustered using DBSCAN (Ester et al., 1996). We look at each cluster’s messages and their corresponding observations to extract any patterns and semantics captured. As shown in Figure 6, we observe a cluster of messages for observations when the goal is visible and another one when another agent is visible. Two clusters correspond to agents seeing neither the goal nor another agent. Notably, the messages in these clusters can come from different agents in different episodes, demonstrating that agents can indeed communicate symmetrically. The clusters indicate that CACL learns to compress meaningful, global state information in messages, allowing other agents to reasonably learn this semantic information.
@@
Table 2 shows the classification results for the two probing tests. Goal visibility is an easier task and all methods’ messages can be effectively used to determine it. In the more difficult goal location task, all methods perform above chance (20%) but CACL’s protocol significantly outperforms baselines. Contrastive learning across different agents’ messages can enable CACL to learn a more global understanding of location from their egocentric viewpoint. We further compare the similarity between messages sent when the goal is at location $[1,1]$ and when the goal is at other coordinates along the diagonal in Appendix A.7, which corroborates with the Goal Location Test in CACL’s capability in encoding global information better. By encoding the goal’s spatial information, CACL agents are more likely able to move directly towards it, and reduce episode length. If other methods simply communicate that a goal is found, agents know to alter their search but are not as precise in direction. This explains why AEComm, PL, and DIAL perform better than IAC but worse than CACL, which also learns much quicker as shown in Figure 2. For completeness, we also provide similar classification results with a one-layer (linear) probe in Appendix A.8.
@@ 7 Conclusion and Future Work
This work introduces an alternative perspective for learning to communicate in decentralized MARL based on the relationship between sent and received messages within a trajectory. Drawing inspiration from multi-view learning, we ground communication using contrastive learning by considering agents’ messages to be encoded views of the same state. We empirically show that our method leads to better performance through a more consistent, common protocol and learns to communicate more global state information. We believe this work solidifies contrastive learning as an effective perspective for learning to communicate and hope it invigorates research into contrastive methods for communication with a focus on consistency. Furthermore, by establishing the connection between multi-view SSL, which has traditionally focused on images, and communication in MARL, we hope to encourage more cross-domain research. Finally, we see contrastive learning as a potential method for simulating human language evolution, and hope to inspire research in this direction.
@@ Appendix A Appendix
A.1 Reinforcement Learning Loss
In Equation 2, we use $L_{RL}$ to denote the RL loss. As our investigation is orthogonal to the choice of RL algorithm, this term can be any RL algorithm dependent on which algorithm is being used. In this work, we use Independent Asynchronous Actor Critic (IA2C) as our base algorithm, denoted as IAC in our experiments (Christianos et al., 2020). Given an agent $i\in N$, it has a policy $\pi^{i}_{\phi}$ and value function $V^{i}_{\theta}$, parameterized by parameters $\phi$ and $\theta$ respectively. The policy loss for agent $i$ is defined as:
$L(\phi)=-\log(\pi(a^{i}_{t}|\Omega^{i}(s_{t});\phi)(r_{t}+\gamma V(\Omega^{i}(s_{t+1});\theta)-V(\Omega^{i}(s_{t});\theta)$ (4)
with the value function minimizes:
$L(\theta)=\lVert V(\Omega^{i}(s_{t});\theta)-(r_{t}+\gamma V(\Omega^{i}(s_{t+1});\theta))\rVert^{2}$ (5)
These two losses are denoted as $L_{RL}$ due to space limit. In the case with communication, the local observation of an agent is simply augmented with received messages.
A.2 Environment Details
@@
Figure 7: Visual illustration of the environments used. Left: Predator-Prey, taken from Koul (2019). Middle: Find-Goal, taken from Lin et al. (2021). Right: Traffic-Junction, taken from Singh et al. (2018)
A.2.1 Predator-Prey
We modify the Predator-Prey implementation by Koul (2019). Our Predator-Prey has a higher communication and coordination requirement than the original Predator-Prey environment. Specifically, for a prey to be captured, it has to be entirely surrounded (i.e. the prey cannot move to another grid position in any actions).
@@ A.2.1 Predator-Prey
A.2.2 Find-Goal
We use the Find-Goal environment implementation provided by Lin et al. (2021). The agents have the goal to find where the goal is in a 15x15 grid world with obstacles.
Unlike in Lin et al. (2021), each agent has a 3x3 field of view (instead of 7x7) to make the task more difficult. Each agent receives a reward of 1 for reaching a goal and an additional reward of 5 if all agents reach the goal. We use a step penalty of -0.01 and an obstacle density of 0.15.
@@ A.2.2 Find-Goal
A.2.3 Traffic-Junction
We use the Traffic-Junction environment implementation provided by Singh et al. (2018). The gridworld is 7x7 with 1 traffic junction. The rate of cars being added has a minimum and maximum of 0.1 and 0.3. We use the easy version with two arrival points and 5 agents. Agents are heavily penalized if a collision happens and have only two actions, namely gas and brake.
@@ A.2.3 Traffic-Junction
A.3 Architecture and Hyperparameters
Figure 8: Architectural illustration for algorithms with communication. To remove communication, the message head is disabled. Grounding module is only relevant to CACL and AEComm. The former is a loss function and the latter is a decoder to reconstruct the encoded observation. Green boxes denote components where gradients of $L_{RL}$ are applied to to. Red boxes denote components where gradients of $L_{CACL}$ and $L_{AEComm}$ are applied to. $L_{AEComm}$ refers to the loss function of AEComm (Lin et al., 2021).
Figure 8 illustrates the components of the architecture used in this work, similar to (Lin et al., 2021). A message head is only used for algorithms with communication, namely CACL, AEComm, PL and DIAL. The Grounding Module refers to mechanisms to ground the messages produced by the message head, used in CACL and AEComm. Colored boxes denote components where gradients are applied to for a particular loss function. In the case of DIAL, gradients would flow from another agent to the agent that sends a message through the message head and message encoder. Unless specified otherwise, we fix all hidden layers to be a size of 32.
We experimented with using the output of the GRU, or hidden state, to condition the message head. Empirically we found that directly conditioning on the observation encoding, as in Lin et al. (2021), led to more stable learning dynamics.
The observation encoder output values of size 32. For Predator-Prey and Traffic Junction, a one-layer fully-connected neural network is used as observation encoder. For Find-Goal, same as Lin et al. (2021), we use a two-layer convolutional neural network followed by a 3-layer fully-connected neural network.
For the message encoder, it outputs values of size 8 in Predator-Prey and Find-Goal with one hidden layer. It outputs values of size 16 in Traffic-Junction with two hidden layers. These configurations are selected based on the best performance of the baseline communication learning algorithm used - DIAL. Messages received are concatenated before passing to message encoders. For all the methods with communication, they produce messages of length 4 ($D=4$) with a sigmoid function as activation. All models are trained with the Adam optimizer (Kingma & Ba, 2014).
@@ A.3 Architecture and Hyperparameters
A.4 CACL Contrastive Loss Curve
Figure 9: CACL’s contrastive loss over time during training for Predator-Prey
Figure 9 shows the CACL’s contrastive loss over time during training for Predator-Prey. This illustrates training convergence of the loss and improved separation of positive and negative samples between the start and end of training.
A.5 Predator-Prey Capture Rate
@@
A.6 Positive Listening
This section describes the loss function we implemented for positive listening, based on Eccles et al. (2019). Given two policies $\pi^{i}$ and $\overline{\pi^{i}}$ of agent $i$ where the latter is the policy with messages zeroed out in the observations, and a trajectory $\tau$ of length $T$, the positive listening loss is written as:
@@ A.6 Positive Listening
A.7 Message similarity for different goal locations in Find-Goal
Figure 10: In Find-Goal, average pairwise cosine similarity between messages sent when the goal is at $[1,1]$ and when the goal is in one of these coordinates $[5,5],[9,9],[13,13]$. Similar to the representation probing test in section 5.6, we generate rollouts from the corresponding algorithms and filter for messages with the goal visible.
To further support that CACL can encode global information better, we measure the message similarity between messages sent when the goal is in $[1,1]$ and when the goal is in other coordinates along the diagonal of the gridworld. As shown in Figure 10, CACL shows a clear downward trend in similarity as the distance from $[1,1]$ increases unlike AEComm. This means CACL learns to encode global locations differently despite learning only with egocentric views. This cooroborates with the Goal Location Test in Table 2 that CACL is able to encode global information better.
A.8 Protocol Representation Probing: 1-Layer
@@
A.9 Zero-shot Cross-play
@@
An advanced form of coordination is working with partners you have not seen during training (Hu et al., 2020). Previous work has focused on coordination through actions (Carroll et al., 2019; Lupu et al., 2021) or pre-test grounding with a common dataset (Gupta et al., 2021) but to our knowledge, no previous work has succeeded in learning a linguistic communication protocol that is robust to zero-shot partners. To assess this advanced robustness, we take trained agents from different methods and random seeds and evaluate them with each other (i.e., zero-shot cross-play) in Predator-Prey and Find-Goal. Given two communication learning methods, $m_{1}$ and $m_{2}$, we sample two agents from each method for Predator-Prey and for Find-Goal, we average over sampling two agents from one method and one agent from the other and vice-versa. For intra-method cross-play, $m_{1}=m_{2}$, we evaluate agents that were trained with the same method but from different random seeds, so they have not been trained with each other. For inter-method cross-play, $m_{1}\neq m_{2}$, we sample agents from two different methods and pair them with each other. Each pairing is evaluated for 10 random seeds each with 10 evaluation episodes. Given that agents are trained in self-play (Tesauro, 1994) without regard for cross-play, we expect severe performance dips.
@@
A.10 Broader Impact
More multi-agent learning systems will be deployed in the real-world as further progress is made in fields of multi-agent learning like MARL. We expect communication to play an essential role in these systems given how real-world problems are inherently complex and partially observable in most cases. Our focus on the decentralized communication setting contributes to the capability of learning more effective and consistent communication protocols. Having more consistent protocols improve mutual intelligibility and pave the way to multi-agent systems in which agents can communicate with unseen agents or even humans.